### v1

# Real-time Surgical Instrument Segmentation in Video Using Point Tracking and Segment Anything

Abstract
The Segment Anything Model (SAM) is a powerful vision foundation model that is revolutionizing the traditional paradigm of segmentation. Despite this, a reliance on prompting each frame and large computational cost limit its usage in robotically assisted surgery. Applications, such as augmented reality guidance, require little user intervention along with efficient inference to be usable clinically. In this study, we address these limitations by adopting lightweight SAM variants to meet the speed requirement and employing fine-tuning techniques to enhance their generalization in surgical scenes. Recent advancements in Tracking Any Point (TAP) have shown promising results in both accuracy and efficiency, particularly when points are occluded or leave the field of view. Inspired by this progress, we present a novel framework that combines an online point tracker with a lightweight SAM model that is fine-tuned for surgical instrument segmentation. Sparse points within the region of interest are tracked and used to prompt SAM throughout the video sequence, providing temporal consistency. The quantitative results surpass the state-of-the-art semi-supervised video object segmentation method on the EndoVis 2015 dataset, with an over 25 FPS inference speed running on a single GeForce RTX 4060 GPU.
Keywords:
Tracking Any Point Segment Anything Surgical Instrument Segmentation.
1 Introduction
Surgical instrument segmentation (SIS) is a fundamental task that provides essential visual cues for various downstream applications of robotic surgery, including augmented reality [1] and surgical scene understanding [2, 3]. Segmenting surgical tools from the tissue background is challenging due to occlusion, blood, smoke, motion artifacts, and changing illumination. While deep learning-based segmentation methods have made significant strides in recent years, achieving high accuracy relies on training using large-scale datasets with annotated images. In surgical computer vision, high-quality annotation is particularly scarce due to the time-consuming, labor-intensive, and expertise-demanding process of labeling.
Recently, the Segment Anything Model (SAM) [4], the first promptable foundation model for image segmentation, has attracted widespread attention. SAM’s demonstrated impressive zero-shot generalization capability along with its flexible prompting framework make it especially useful for enabling downstream applications. In surgical scenarios, however, the application of SAM faces two challenges. Firstly, the huge computational cost of its heavyweight image encoder architecture, especially when processing high-resolution images, hinders its real-time inference capabilities [5]. Furthermore, numerous studies have reported significant performance degradation of SAM on medical images [20, 21], including images in surgical scenes [19]. In this study, we adopt the lightweight SAM variant to facilitate inference efficiency. Furthermore, we investigate the point prompt-based fine-tuning strategy for MobileSAM [7] to mitigate the performance degradation associated with the lightweight network architecture.
Despite SAM’s strong automatic mask generation ability, achieving expected segmentation results in practice often requires appropriate prompts. Providing specific points or descriptive text of the target object can significantly improve the segmentation accuracy. Leveraging the long-term tracking capabilities of the Tracking Any Point (TAP) models, we employ an online point tracker, CoTracker [9], to provide sparse point prompts for SAM. Similar to SAM-PT [8] and DEVA [23], our pipeline decouples video object segmentation (VOS) into image-level segmentation, which can be task-specific, and a universal temporal propagation. Compared to end-to-end VOS, our “tracking-by-detection” framework can take full advantage of smaller image-level datasets via fine-tuning a task-specific image segmentation model and using it in tandem with a point tracker to maintain temporal consistency.
In summary, our contribution is threefold: (1) we present a real-time video surgical instrument segmentation framework that achieves superior segmentation performance and is suitable for clinical usage due to its good efficiency; (2) we investigate the point prompt-based fine-tuning strategy (will open source) for lightweight SAM using surgical datasets, and the model fine-tuned on only two datasets shows promising generalization on unseen surgical videos; (3) a naive combination of SAM and TAP cannot achieve ideal performance from both the segmentation accuracy and the inference efficiency perspectives for online SIS, while our presented pipeline significantly enhances the performance on both fronts.
2 Related Work
2.0.1 Tracking Any Point:
The goal of Tracking Any Point (TAP) is to estimate the motion of arbitrary physical points throughout a video. TAP-Vid [12] first formalized this task alongside a benchmark dataset and baseline method for TAP. Recent work has showcased the promising online TAP capability and exhibited great robustness to occlusion and exit from the field of view. PIPs++ [13] and TAPIR [14] demonstrate a substantial robustness under occlusion and achieve real-time inference speed on high-resolution video. Notably, CoTracker [9] achieves state-of-the-art tracking performance by jointly tracking a set of query points. CoTracker is an online algorithm that processes video sequentially through a sliding window. Optical flow [10, 11] can be used for TAP but tends to accumulate errors over time and faces challenges in handling occlusions, which are common occurrences in surgical scenarios.
2.0.2 Segment Anything:
SAM is the first vision foundation model for image segmentation, trained over the SA-1B dataset consisting of 1 billion high-quality annotated images [4]. SAM demonstrates impressive zero-shot inference capability on natural images and supports flexible prompts. Nevertheless, SAM’s performance often declines in specific fields [19, 20, 21, 22], which can be attributed to a substantial domain gap. Much research has been dedicated to adapting SAM to medical images [15, 18] including surgical images [16, 17]. SurgicalSAM [16] and AdaptiveSAM [17] are adapted to the surgical domain by providing class and text prompts. However, neither has real-time inference speed. The computational cost of SAM stems from its heavy image encoder, with some research [5, 6, 7, 34] aiming to accelerate SAM’s inference and reduce the demand for computation resources.
3 Methods
Our proposed framework consists of two key components: a point tracker and a point-based segmentation model. Both components can be flexibly replaced with state-of-the-art models. As shown in Fig. 1, the pipeline can be described as follows. To begin with, the first frame mask of the video sequence is generated to indicate the region of interest (ROI) where query points are initialized. Subsequently, a set of query points is selected within the ROI based on a sampling strategy. After this pre-processing, we employ a point tracker to track these query points and utilize them as prompts at each frame for the segmentation model.
In Section 3.1, we illustrate the pre-processing that is used to initialize query points. In Section 3.2, we formalize and clarify the proposed TAP + SAM framework. In Section 3.3, we introduce the SAM fine-tuning strategy employed in this study.
Figure 1: The pre-processing workflow to generate the query points. The segmentation model can be semi-automatic, i.e., SAM + bounding box prompt (bottom row), or fully automatic, i.e., CLIPSeg + fixed text prompt “surgical tool” (top row). Note that, without the initial mask, just manually picking query points is also feasible.
3.1 Pre-processing
We use the SAM model to generate the initial mask by simply inputting a few points or bounding boxes. Inspired by recent advancements in vision-language models, several zero-shot segmentation models based on text prompts have been developed [24, 25]. To achieve the “fully” automatic pipeline, we incorporate a text-promptable segmentation model CLIPSeg [24] to automatically generate the initial mask by setting a text prompt, “surgical tool”. While CLIPSeg can only provide a coarse initial mask, it remains feasible for query point selection as long as its output roughly covers the region of the target instrument.
We initialize query points using the first frame mask. We investigate various query point sampling strategies, including random sampling, uniform sampling on a grid, SIFT keypoints, Shi-Tomasi corner points, and K-Medoids [35] clustering centers. Further details on these sampling strategies are provided in the Supplementary Material. We selected K-Medoids clustering centers because they ensure even partitioning of the entire cluster. The number of medoids assigned to each instance ranges from 1 to 9, and we choose 5 in the experiments. Let
$P_{0}=\{(p_{i},t_{0})\},\,\,\,\,p_{i}=(x_{i},y_{i}),\,\,i=1,...,N$ (1)
be the initial set of query point locations $p_{i}$ at time $t_{0}$.
Figure 2: The overview of our video surgical instrument segmentation pipeline that combines a segmentation model $Seg(\cdot)$ and a point tracker $TAP(\cdot)$.
3.2 Tracking Any Point + Segment Anything
We describe our framework with reference to Fig. 2. Given a video $V=\{I_{t}\},$ in which $I_{t}\in R^{2}$ is the image at time $t$, along with a set of initial query points $P_{0}$ in $I_{0}$, we use TAP to predict the query points $P_{t}=\{(p_{i},\;t)\}$, at time $t$
$P_{t}=TAP(V,\;P_{0}),$ (2)
and we use $P_{t}$ and the current image $I_{t}$ in a segmentation model
$M_{t}=Seg(I_{t},P_{t}).$ (3)
to produce the mask $M_{t}$.
We incorporate the state-of-the-art model, CoTracker [9], as the off-the-shelf online point tracker to propagate the initial query points throughout the video sequence. CoTracker takes a short video clip consisting of several frames as the input. It processes video frames in a serial fashion via a 4-frame sliding window, making it suitable for online applications. We also integrated PIPs++ [13] and TAPIR [14] into our software, but only CoTracker is used for experiments in this paper. For the segmentation model $Seg$, we adopt the fine-tuned MobileSAM to enable real-time processing throughout the entire pipeline, while achieving accurate segmentation.
3.3 Fine-tuning the Segment Anything Model
As for the segmentation model, we first tested two state-of-the-art lightweight SAM variants, MobileSAM and Light HQ-SAM [6]. However, we observed that both methods perform poorly in situations where specularity, blood, or weak lighting is present, as shown in Fig. 4. Driven by this limitation, enhancing the generalization of the lightweight SAM for surgical instrument segmentation becomes crucial. The state-of-the-art MedSAM [15] has demonstrated that fully fine-tuning SAM for medical images can yield promising results. Fully fine-tuning refers to freezing the prompt encoder and updating both the image encoder and mask decoder. Compared with strategies that only update the mask decoder or introducing the adapter layer, fully fine-tuning achieves superior performance.
Figure 3: The pipeline of SAM fine-tuning using points. The input for the SAM model consists of images paired with points within the target area. The green rectangle represents the MobileSAM architecture.
In this study, we investigate the fully fine-tuning strategy for the most widely used lightweight SAM variant, MobileSAM, to optimize its capability for surgical instrument segmentation. As depicted in Fig. 3, the network architecture of MobileSAM is consistent with the original SAM. Unlike using bounding box prompts in MedSAM, we utilize point prompts during training to maintain consistency with the prompt type used during inference. Both the image encoder and mask decoder are learnable and updated during the fine-tuning. Due to the small number of learnable parameters of MobileSAM (10.13M), the cost of computation is significantly reduced. We train our model on 4 V100 GPUs for 50 epochs. The training is based on the pre-trained MobileSAM weight. In datasets with instance-level labels, we randomly sample 5 points as prompts within the area of each instance. For datasets with binary labels, we randomly sample 5 points as prompts within the segmented region. The loss $L$ utilizes an unweighted combination of binary cross entropy loss and Dice loss [27], represented as
$L=L_{BCE}\;+\;L_{Dice}$ (4)
We used the AdamW [32] optimizer for training, with a batch size of 32. The initial learning rate is set to 1e-5 and follows a cosine decay schedule. All images are resized to 1024 by 1024 and undergo random up-down and left-right flip data augmentation.
4 Experiments and Results
4.0.1 Dataset:
We conduct quantitative comparisons on the EndoVis 2015 [30] and the UCL dVRK [28] datasets to validate the feasibility and performance of our proposed framework. Additionally, we display qualitative results on these two datasets as well as results on the unlabeled high-resolution STIR [29] dataset. The EndoVis 2015 dataset provides annotations for both rigid and articulated robotic instruments in laparoscopic surgery. In this study, we focus solely on the robotic instruments. For both datasets, we keep the original data split for fine-tuning and testing. The frame rates of the videos and annotations in the EndoVis 2015 is 25 FPS. In the UCL dVRK dataset, images are collected and annotated at 6.67 FPS. In the STIR dataset, the video frame rate is 30 FPS.
4.0.2 Quantitative Results:
We perform extensive comparisons between different SAM variants and the state-of-the-art semi-supervised VOS method, XMem [31]. We adopt the standard VOS metric [33], $\mathcal{J}\&\mathcal{F}$, the average of Jaccard region similarity ($\mathcal{J}$) and contour accuracy ($\mathcal{F}$) for quantitative comparison. Table 1 and 2 show the quantitative results of the EndoVis 2015 and UCL dVRK datasets, respectively. $\mathcal{J}$ and $\mathcal{F}$ results are provided in the Supplementary Material. The inference speed of our fine-tuned MobileSAM is about 40 milliseconds, while the ViT-H SAM’s inference speed is about 0.9 seconds. The CoTracker’s frame rate fluctuates in the 50-60 FPS range. All experiments are running on a single GeForce RTX 4060 GPU.
Table 1: Quantitative Results ($\mathcal{J}\&\mathcal{F}$) on the EndoVis 2015 Dataset. The abbreviations PT and FF-SAM stand for point tracker and fully fine-tuned MobileSAM, respectively. The same abbreviations are used in Table 2.
Method Video 1 Video 2 Video 3 Video 4 Video 5 Video 6
PT+MobileSAM [7] 79.9 87.8 84.7 59.5 65.3 74.9
PT+Light HQ-SAM [6] 81.3 88.7 80.1 58.8 63.5 75.1
PT+ViT-H SAM [4] 87.2 89.3 86.9 81.2 77.0 81.9
PT+FF-SAM 89.0 92.1 88.9 84.6 86.3 85.5
XMem [31] 87.3 91.7 88.7 84.2 83.1 83.8
Table 2: Quantitative Results ($\mathcal{J}\&\mathcal{F}$) on the UCL dVRK Dataset.
Method Video 11 Video 12 Video 13 Video 14
PT+MobileSAM [7] 55.6 33.6 50.6 64.3
PT+Light HQ-SAM [6] 36.6 70.6 60.7 80.2
PT+ViT-H SAM [4] 73.6 81.2 79.6 80.8
PT+FF-SAM 89.9 89.5 94.0 91.1
XMem [31] 93.5 93.4 94.4 93.1
Figure 4: Visualization of qualitative results from our proposed methods on different datasets. The point tracker is CoTracker, which we pair with different SAM variants. The red dots are the query points for prompting SAM. The images are from STIR, EndoVis 2015, and UCL dVRK, respectively (top to bottom).
4.0.3 Qualitative Results:
Fig. 4 displays a few qualitative results of various methods on different datasets. To demonstrate the generalization of the proposed method, we test our framework, which is fine-tuned using the EndoVis 15 and UCL dVRK dataset, on some videos from the STIR dataset. Note that the model has never seen this dataset.
5 Discussion
Overall, our proposed method outperforms the state-of-the-art semi-supervised VOS model, XMem, on the EndoVis 2015 dataset. Notably, in the one-minute long videos, Video 5 and 6, our proposed method exhibits obvious improvement compared to XMem. This shows that our framework can robustly leverage the temporal information provided by universal TAP methods. The significance of fine-tuning is demonstrated by the substantial improvement compared to non-fine-tuned SAMs, even that of the powerful ViT-H SAM. Note that the inference speed of the ViT-H SAM is far away from real-time. In the UCL dVRK dataset, our proposed method also achieves promising results, although it does not outperform XMem. All four UCL dVRK testing videos are captured under very weak illumination conditions, which is unrealistic in real surgery. The dark scenes make it challenging for SAM to distinguish the boundary from the tissue background, especially when there are no such weak lighting conditions in the training set. In contrast, the mask propagation-based XMem takes the full first frame ground truth for its inference. The low video frame rate (6.67 FPS) of the UCL dVRK dataset also poses an obstacle for the TAP, thereby making it hard to provide effective point prompts throughout the video sequence. Compared to XMem, our method has other advantages. XMem requires an accurate first frame mask, while our pipeline only needs a known text prompt “surgical tool”. Furthermore, XMem cannot recognize new objects during the video, while our method can tackle this by enabling users to easily pick a few new query points, instead of providing a high-quality mask. The quantitative results and qualitative examples depicted in Fig. 4 show the superiority of the fine-tuned MobileSAM over other lightweight SAM variants and even the ViT-H SAM.
6 Conclusion
In this study, we present a novel framework using a universal TAP and a fine-tuned lightweight SAM for real-time surgical instrument segmentation in video. Its commendable efficiency and accuracy makes it suitable for applications in clinical settings. We investigate the availability of fine-tuning MobileSAM using point prompts and demonstrate the importance of fine-tuning for SIS. Extensive experiments validate the advancement of our proposed pipeline. Furthermore, our SAM + TAP pipeline demonstrates the potential to serve as a strong VOS baseline by integrating other image segmentation models.